The measure of an angle is 44 more than the measure of its supplement. Find the measure of the angles.
Thanks!

Answer by sdmmadam@yahoo.com(530)

Let the first angle be A degrees and
let the second angle be B degrees
Given A is 44 more than B
Therefore A = B + 44
Also B is the supplement of A
Therefore A + B = 180
(B + 44) + B = 180 (putting the value of A in terms of B by data)
By additive associativity and commutativity
2B + 44 = 180
2B = 180 - 44 ( by transposing, change side then change sign)
2B = 136
B = 136/2 = 68 and therefore angle B = 68 degrees
And A = B + 44 = 68 + 44 = 112 and therefore angle A = 112 degrees.
Answer : The two angles are 112 and 48 degrees respectively.
Verification: Given that the two angles are supplementary.
Now A + B = 112 + 68 = 180 which is correct.